# Three-dimensional Abelian and nonAbelian gauge Higgs theories

### v1

@@ 9.2.2 O(5) multicritical behavior of the two-flavor SU(2) gauge Higgs model
The three transition lines that occur for $N_{c}=2$ intersect on the $v=0$ axis, see Fig. 19 (we discuss the behavior at fixed $\gamma$; in the full phase diagram, we have a $\gamma$ dependent line of multicritical points). The behavior at the intersection point depends on $N_{f}$. For $N_{f}>2$, the intersection point with symmetry Sp($N_{f}$)/$\mathbb{Z}_{2}$ is expected to be a first-order transition point, as predicted by an analysis based on the corresponding multicritical LGW theory [97, 167]. [^45] We should also mention that the field-theory model with $v=0$ admits a charged FP for large values of $N_{f}$. Thus, in this case a charged multicritical behavior can also appear. Therefore, we expect the three transition lines to be of first order close to the intersection. More precisely, the discussion presented in Sec. 9.1.2 implies that for $N_{f}<N_{f}^{*}(3)$ the transitions are of first order for any value of $v$. On the other hand, for $N_{f}>N_{f}^{*}(3)$, charged continuous transitions are possible for $v>0$. Even in this case, we however expect first-order transitions for any $v<0$, and also for sufficiently small positive values of $v$.
@@
As discussed in A.4, the multicritical behavior is controlled by two relevant RG perturbations, represented by two scaling fields $u_{1}$ and $u_{2}$ that are functions of the model parameters and have positive RG dimensions $y_{1}$ and $y_{2}$. In the case of the O(5)-symmetric MCP we can identify $u_{1}\approx v$ and $u_{2}\approx 1-\beta/\beta_{\rm mc}$, whose RG dimensions are determined by the O(5) vector universality class [173, 176]. They are given by $y_{1}=y_{2,2}=1.832(8)$ [173], where $y_{2,2}$ is the RG scaling dimension of the coupling associated to the spin-two perturbation of the O(5)-vector FP, and $y_{2}=1/\nu_{{\rm O}(5)}=1.2818(10)$ (using the best available estimate [294] $\nu_{{\rm O}(5)}=0.7802(6)$ for the length-scale exponent of the O(5) vector universality class).
@@
9.2.3 Phase diagram of multicomponent SU($N_{c}$) gauge Higgs models with $N_{c}\geq 3$
Figure 20: Sketches of the phase diagram at fixed $\gamma\geq 0$ expected for $N_{c}\geq 3$, and $N_{f}<N_{c}$ (left), $N_{f}=N_{c}$ (middle), $N_{f}>N_{c}$ (right).
@@ 9.2.3 Phase diagram of multicomponent SU($N_{c}$) gauge Higgs models with $N_{c}\geq 3$
9.3 Critical behaviors described by the charged fixed point of the SU($N_{c}$) gauge Higgs field theory
Figure 21: FSS behavior of the 3D SU(2) gauge LNAH model along the disorder-Higgs transition. Results for $N_{f}=40$, $v=1$, and $\gamma=1$, varying the inverse-temperature $\beta$. The left panel shows data for $\xi_{P}/L$, where $\xi_{P}$ is the second-moment correlation length computed from the correlation function $G_{P}$ defined in Eq. (105), versus $X=(\beta-\beta_{c})L^{1/\nu}$, using the best estimates $\beta_{c}=1.18625$ and $\nu=0.745$. The excellent collapse of the data confirms that the transition is continuous. The inset shows the same data versus $\beta$. The right panel shows the rescaled susceptibility $\chi_{P}$ associated with $G_{P}$, which asymptotically behaves as $\chi_{P}\approx L^{2-\eta_{P}}{\cal X}_{P}(\xi_{P}/L)$, see Eq. (153), with $\eta_{P}=0.87(1)$.
An important issue is the relation between the statistical gauge model (99) and the corresponding NAHFT (85), which is the field theory with the same field content and the same local and global symmetries. In particular, one would like to identify the continuous transitions that can be described by the stable charged FP of the RG flow of the continuum SU($N_{c}$) gauge Higgs field theory, see Sec. 8.1.3. The analogous problem for charged transitions in AH models has been discussed in Sections 5 and 6.
@@ 9.3 Critical behaviors described by the charged fixed point of the SU($N_{c}$) gauge Higgs field theory
Refs. [392, 99] report a numerical study of the SU(2) gauge model ($N_{c}=2$) aiming at identifying possible charged transitions for large values of $N_{f}$. Continuous transitions are observed for $N_{f}=30,40$, and 60, keeping $v=1$ and $\gamma=1$ fixed and varying $\beta$. Some numerical FSS results are shown in Fig. 21 for $N_{f}=40$; they clearly show that the transition is continuous. To confirm that these large-$N_{f}$ critical behaviors are associated with the stable charged FP of the NAHFT, one may compare the numerical estimates of the correlation-length exponent $\nu$ with the field-theoretical 1/$N_{f}$ result reported in Eq. (93). Such a comparison is reported in Fig. 22, which shows that numerical data and large-$N_{f}$ field-theoretical predictions are in nice agreement. In particular, the estimate $\nu=0.745(10)$ for $N_{f}=40$ appears in reasonable agreement (the small difference can be interpreted as a 1/$N^{2}_{f}$ correction, see Ref. [392]) with the value $\nu\approx 0.757$ obtained from the $O(N_{f}^{-1})$ formula reported in Eq. (93). The results of Refs. [392, 99] allow one to estimate $N^{*}_{f}$ in three dimensions for $N_{c}=2$. Since only first-order transitions are observed for $N_{f}=20$ for different values of $v$ and $\gamma$, and continuous transitions are found for $N_{f}=30$, one may conclude that $N_{f}^{*}=25(4)$ [392]. Note that this value is significantly smaller than $N_{f}^{*}\approx 375$ close to four dimensions. This difference is not surprising, since a similar difference was observed for the AHFT, see Sec. 4.1.2.
Figure 22: (Adapted from Ref. [392]) Numerical estimates of the critical exponent $\nu$ versus $1/N_{f}$ for the SU(2) gauge LNAH model (99). For comparison we also report the $O(1/N_{f})$ theoretical prediction, Eq. (93) (solid line), and a next-to-leading interpolation including a $1/N_{f}^{2}$ term, i.e. $\nu=1+a_{1}/N_{f}+a_{2}/N_{f}^{2}$ (dashed line), with $a_{2}\approx-30$.
We also mention the numerical results reported in Ref. [99], which show that the charged critical behavior for $N_{f}=40$ and $\gamma=1$ definitely differs from that observed for $\gamma\to\infty$, confirming that gauge fluctuations are relevant along the large-$\gamma$ charged transition line.
The agreement between the NAHFT results and the lattice estimates allows us to conjecture that the continuous transitions observed in the lattice model can be associated with the stable charged FP of the RG flow of the SU(2) gauge NAHFT formally defined by the Lagrangian (85) and its corresponding functional path integral. We expect that analogous results hold for $N_{c}\geq 3$, although larger values of $N_{f}$ are likely required to observe charged continuous transitions. We finally remark that the existence of these 3D charged universality classes implies the existence of a well-defined nonperturbative continuum limit of the SU($N_{c}$) gauge NAHFTs, for a sufficiently large number of scalar components.
@@ 9.4.1 Lattice Hamiltonian
The fundamental variables are real matrices $\Phi^{af}_{\bm{x}}$, with $a=1,...,N_{c}^{2}-1$ (color index) and $f=1,...,N_{f}$ (flavor index), defined on the lattice sites, and gauge fields $U_{{\bm{x}},\mu}\in{\rm SU}(N_{c})$ associated with the lattice links [100]. The Hamiltonian and the corresponding partition function are [259, 98, 377, 383, 384, 385]
@@
where the matrix $\widetilde{U}^{ab}_{{\bm{x}},\mu}$ is the adjoint representation of the original link variable $U_{{\bm{x}},\mu}$. The matrix $\widetilde{U}^{ab}_{{\bm{x}},\mu}$ is explicitly given by $\widetilde{U}^{ab}_{{\bm{x}},\mu}=2\,{\rm Tr}(\,U^{\dagger}_{{\bm{x}},\mu}T^{a}U_{{\bm{x}},\mu}T^{b}\,)$, where $a,b=1,...,N_{c}^{2}-1$, and $T^{a}$ are the $N_{c}^{2}-1$ generators in the fundamental representation, with normalization ${\rm Tr}\,T^{a}T^{b}=\frac{1}{2}\delta^{ab}$. In the following we fix $J=1$, so that energies are measured in units of $J$. The scalar potential term $H_{V}$ is written as
@@
$U_{{\bm{x}},\mu}\to V_{\bm{x}}U_{{\bm{x}},\mu}V_{{\bm{x}}+\hat{\mu}}^{\dagger}\qquad\Phi^{af}_{\bm{x}}\to\sum_{b}\widetilde{V}_{\bm{x}}^{ab}\Phi^{bf}_{\bm{x}},$ (111)
where $V_{\bm{x}}$ is an SU($N_{c}$) matrix and $\widetilde{V}_{\bm{x}}$ is the corresponding matrix in the adjoint representation. The Hamiltonian is also invariant under the transformations $U_{{\bm{x}},\mu}\to z(x_{\mu})U_{{\bm{x}},\mu}$, where $z(x_{\mu})$ is an element of the gauge-group center ${\mathbb{Z}}_{N_{c}}$ that depends only on $x_{\mu}$ (the $\mu$-th component of the position vector). When this symmetry is not spontaneously broken, Wilson loops obey the area law and color charges transforming in the fundamental representation are confined.
One may again consider a simplified model, requiring the scalar fields to obey the fixed length constraint ${\rm Tr}\,\Phi_{\bm{x}}^{t}\Phi_{\bm{x}}=2$. Formally, this model can be obtained by taking the limit $u,r\to\infty$ keeping the ratio $r/u=-2$ fixed. The corresponding lattice Hamiltonian reads
@@
The critical properties of the scalar fields can be monitored by using the correlation functions $G_{P}$ of the gauge-invariant bilinear operator
@@ 9.4.2 Higgs phases and phase transitions
For $N_{f}\geq 2$ the lattice models show different Higgs phases associated with different symmetry-breaking patterns. The symmetries of the Higgs phases can be inferred by performing a mean-field analysis [96, 259, 98]—critical fluctuations are only relevant along the transition lines. The symmetries of the Higgs phases are indeed the symmetries of the minima of the local scalar potential $V(\Phi)={v\over 4}\,{\rm Tr}\,(\Phi^{t}\Phi)^{2}$ in the fixed-length limit ${\rm Tr}\,\Phi^{t}\Phi=2$. In the following we summarize the main properties of these phases, which depend on the number of colors $N_{c}$, of flavors $N_{f}$, and on the parameter $v$ [96, 259, 98]. Moreover, their nature also depends on the behavior of the gauge modes, which may undergo topological transitions due to the fluctuations of variables associated with the gauge-group center ${\mathbb{Z}}_{N_{c}}$. Numerical studies for $N_{c}=2$ and $N_{c}=3$ have been reported in Refs. [259, 98, 407].
@@
Two qualitatively different phase diagrams are found, depending on the number of colors $N_{c}$ and of flavors $N_{f}$. For $N_{f}\leq N_{c}^{2}-1$, independently of the specific value of $N_{c}$, there is a single Higgs phase that lies in the half-space $v<0$ for large values of $\beta$. Indeed, the mean-field analysis shows that, for $\beta\to\infty$ the model orders only for $v<0$. For $v>0$, the system is disordered for $\beta\to\infty$.
[…]
@@
The possible existence of topological transitions associated with the gauge-group center ${\mathbb{Z}}_{N_{c}}$ can also be inferred by looking at the behavior of the model in the limit $\beta\to\infty$ keeping $\kappa\equiv\beta\gamma$ fixed, see Ref. [98] for details. For $v>0$ the relevant configurations, i.e., those that minimize the scalar potential and the scalar kinetic energy $H_{K}$, are characterized by $\widetilde{U}_{{\bm{x}},\mu}=1$, so $U_{{\bm{x}},\mu}=\lambda_{{\bm{x}},\mu}\in{\mathbb{Z}}_{N_{c}}$. Therefore, in this limit the model (112) reduces to the lattice ${\mathbb{Z}}_{Q}$ gauge theory with $Q=N_{c}$ discussed in Sec. 3.5.2. In three dimensions, this lattice discrete gauge model undergoes a transition at a finite value of the gauge coupling $K$. For example, the case $N_{c}=2$ corresponds to a lattice ${\mathbb{Z}}_{2}$ gauge theory [119], which presents a small-$K$ confined phase and a large-$K$ deconfined phase (which may carry topological order at the quantum level [20]), separated by a critical point at $K_{c}=0.761413292(11)$ [86, 316]. If the ${\mathbb{Z}}_{N_{c}}$ gauge transition persists for finite values of $\beta$, then, when varying $\gamma$ or $\kappa$, we may have different low-temperature Higgs phases with the same global symmetry that differ in the large-scale behavior of the ${\mathbb{Z}}_{N_{c}}$ modes.
@@
These arguments indicate that there are two different positive-$v$ Higgs phases, which differ for the behavior of the topological modes associated with the gauge-group center ${\mathbb{Z}}_{N_{c}}$ modes [96, 259]. Some numerical evidence of the existence of topological ${\mathbb{Z}}_{2}$ transitions in SU($2$) LNAH models with four adjoint scalar fields can be found in Ref. [98].

### v3

@@ 9.2.3 O(5) multicritical behavior of the two-flavor SU(2) gauge Higgs model
We now focus on a particular region of the phase diagram of SU(2) gauge LNAH models with $N_{f}$ flavors. We discuss the nature of the transition point that lies on the $v=0$ axis, where the transition lines intersect, see Fig. 19 (we discuss the behavior at fixed $\gamma$; in the full phase diagram, we have a $\gamma$ dependent line of such intersection points). Note that the global symmetry of the model enlarges to Sp($N_{f}$)/$\mathbb{Z}_{2}$ for $v=0$.
The behavior at the intersection point depends on $N_{f}$. For $N_{f}>2$ and $N_{f}$ not too large, the intersection point is expected to be a first-order transition point, as predicted by an analysis based on the corresponding multicritical LGW theory [98, 198]. Therefore, we expect the three transition lines to be of first order close to the intersection. Note that the arguments presented in Sec. 9.2.1 imply that, for $N_{f}>N_{f}^{*}(3)$, charged continuous transitions are possible for $v>0$ (for $v<0$ only first-order transitions are possible). If the intersection point is a first-order transition, we expect first-order transitions for sufficiently small positive values of $v$ also in this case. Finally, note that the field-theory model with $v=0$ admits a charged FP for large values of $N_{f}$. Thus, for $N_{f}\gg 1$ a charged multicritical behavior can also appear.
@@
As discussed in A.4, the multicritical behavior is controlled by two relevant RG perturbations, represented by two scaling fields $u_{1}$ and $u_{2}$ that are functions of the model parameters and have positive RG dimensions $y_{1}$ and $y_{2}$. In the case of the O(5)-symmetric MCP we can identify $u_{1}\approx v$ and $u_{2}\approx 1-\beta/\beta_{\rm mc}$, whose RG dimensions are determined by the O(5) vector universality class [204, 152]. They are given by $y_{1}=y_{2,2}=1.832(8)$ [204] ($y_{2,2}=1.838(10)$ from Monte Carlo simulations [435]), where $y_{2,2}$ is the RG scaling dimension of the coupling associated to the spin-two perturbation of the O(5)-vector FP, and $y_{2}=1/\nu_{{\rm O}(5)}=1.2818(10)$ (using the best available estimate [322] $\nu_{{\rm O}(5)}=0.7802(6)$ for the length-scale exponent of the O(5) vector universality class).
@@
9.2.4 Lattice SU($N_{c}$) gauge Higgs models with $N_{c}\geq 3$
Figure 20: Sketches of the phase diagrams of 3D SU($N_{c}$) gauge LNAH models with $N_{f}$ scalar fields in the fundamental representation, at fixed $\gamma\geq 0$, for $N_{c}\geq 3$, and $N_{f}<N_{c}$ (left), $N_{f}=N_{c}$ (middle), $N_{f}>N_{c}$ (right).
@@ 9.2.4 Lattice SU($N_{c}$) gauge Higgs models with $N_{c}\geq 3$
9.3 Critical behaviors associated with the charged fixed point of the SU($N_{c}$) gauge Higgs field theory
Figure 21: FSS behavior of the 3D SU(2) gauge LNAH model along the disorder-Higgs transition line. Results for $N_{f}=40$, $v=1$, and $\gamma=1$, varying the inverse temperature $\beta$. The left panel shows data for $\xi_{P}/L$, where $\xi_{P}$ is the second-moment correlation length computed from the correlation function $G_{P}$ defined in Eq. (108), versus $X=(\beta-\beta_{c})L^{1/\nu}$, using the best estimates $\beta_{c}=1.18625$ and $\nu=0.745$. The excellent collapse of the data confirms that the transition is continuous. The inset shows the same data versus $\beta$. The right panel shows the rescaled susceptibility $\chi_{P}$ of the operator $P_{\bm{x}}$, which asymptotically behaves as $\chi_{P}\approx L^{2-\eta_{P}}{\cal X}_{P}(\xi_{P}/L)$, see Eq. (158), with $\eta_{P}=0.87(1)$.
An important issue is the relation between the statistical lattice gauge model (102) and the corresponding NAHFT (88), which is the field theory with the same field content and the same local and global symmetries. In particular, one would like to identify the continuous transitions that can be described by the stable charged FP of the RG flow of the continuum SU($N_{c}$) gauge Higgs field theory, see Sec. 8.1.3. The analogous problem for charged transitions in AH models has been discussed in Secs. 5 and 6.
@@
Refs. [434, 100] report a numerical study of the SU(2) gauge model ($N_{c}=2$) aiming at identifying possible charged transitions for large values of $N_{f}$. Continuous transitions are observed for $N_{f}=30,40$, and 60, keeping $v=1$ and $\gamma=1$ fixed and varying $\beta$. Some numerical FSS results are shown in Fig. 21 for $N_{f}=40$; they clearly show that the transition is continuous. We also mention that the observed FSS behavior for $N_{f}=40$ and $\gamma=1$ differs from that observed for the same value of $N_{f}$ ($N_{f}=40$) in the limit $\gamma\to\infty$ [100], confirming that gauge fluctuations are relevant along the large-$\gamma$ charged transition line.
To confirm that these large-$N_{f}$ transitions have a critical behavior associated with the stable charged FP of the NAHFT, one may compare the numerical estimates of the correlation-length exponent $\nu$ with the field-theoretical 1/$N_{f}$ result reported in Eq. (96). Such a comparison is reported in Fig. 22, which shows that numerical data and large-$N_{f}$ field-theoretical predictions are in nice agreement. In particular, the estimate $\nu=0.745(10)$ for $N_{f}=40$ appears in reasonable agreement (the small difference can be interpreted as a 1/$N^{2}_{f}$ correction, see Ref. [434]) with the value $\nu\approx 0.757$ obtained from the $O(N_{f}^{-1})$ formula reported in Eq. (96). The results of Refs. [434, 100] allow one to estimate $N^{*}_{f}$ in three dimensions for $N_{c}=2$. Since only first-order transitions are observed for $N_{f}=20$ for different values of $v$ and $\gamma$, and continuous transitions are found for $N_{f}=30$, one may conclude that $N_{f}^{*}=25(4)$ [434]. Note that this value is significantly smaller than $N_{f}^{*}\approx 375$, the result in four dimensions. This difference is not surprising, since a similar difference was observed for the AHFT, see Sec. 4.1.2.
Figure 22: (Adapted from Ref. [434]) Numerical estimates of the critical exponent $\nu$ versus $1/N_{f}$ for the SU(2) gauge LNAH model (102). For comparison we also report the $O(1/N_{f})$ theoretical prediction, Eq. (96) (solid line) and a next-to-leading interpolation including a $1/N_{f}^{2}$ term, i.e. $\nu=1+a_{1}/N_{f}+a_{2}/N_{f}^{2}$ (dashed line), with $a_{2}\approx-30$.
The agreement between the NAHFT results and the lattice numerical estimates nicely supports the conjecture that the continuous transitions observed in the lattice model can be associated with the stable charged FP of the RG flow of the SU(2) gauge NAHFT formally defined by the Lagrangian (88) and the corresponding functional path integral. We expect that analogous results hold for $N_{c}\geq 3$, although larger values of $N_{f}$ are likely required to observe charged continuous transitions. We finally remark that the existence of these 3D charged universality classes implies the existence of a well-defined nonperturbative continuum limit of the SU($N_{c}$) gauge NAHFTs, for a sufficiently large number of scalar components.
@@ 9.4.1 Lattice Hamiltonian
The fundamental variables are real matrices $\Phi^{af}_{\bm{x}}$, with $a=1,...,N_{c}^{2}-1$ (color index) and $f=1,...,N_{f}$ (flavor index), defined on the lattice sites, and gauge fields $U_{{\bm{x}},\mu}\in{\rm SU}(N_{c})$ associated with the lattice links [101]. The Hamiltonian and the corresponding partition function are [289, 99, 409, 415, 416, 417, 418]
@@
where the matrix $\widetilde{U}^{ab}_{{\bm{x}},\mu}$ is the adjoint representation of the original link variable $U_{{\bm{x}},\mu}$. The matrix $\widetilde{U}^{ab}_{{\bm{x}},\mu}$ is explicitly given by $\widetilde{U}^{ab}_{{\bm{x}},\mu}=2\,{\rm Tr}(\,U^{\dagger}_{{\bm{x}},\mu}T^{a}U_{{\bm{x}},\mu}T^{b}\,)$, where $a,b=1,...,N_{c}^{2}-1$, and $T^{a}$ are the $N_{c}^{2}-1$ generators in the fundamental representation, with normalization ${\rm Tr}\,T^{a}T^{b}=\frac{1}{2}\delta^{ab}$. In the following we fix $J=1$, so t energies are measured in units of $J$. The scalar potential term $H_{V}$ is written as
@@
$U_{{\bm{x}},\mu}\to V_{\bm{x}}U_{{\bm{x}},\mu}V_{{\bm{x}}+\hat{\mu}}^{\dagger},\qquad\Phi^{af}_{\bm{x}}\to\sum_{b}\widetilde{V}_{\bm{x}}^{ab}\Phi^{bf}_{\bm{x}},$ (114)
where $V_{\bm{x}}$ is a SU($N_{c}$) matrix and $\widetilde{V}_{\bm{x}}$ is the corresponding matrix in the adjoint representation. The Hamiltonian is also invariant under the transformations $U_{{\bm{x}},\mu}\to z(x_{\mu})U_{{\bm{x}},\mu}$, where $z(x_{\mu})$ is an element of the gauge-group center ${\mathbb{Z}}_{N_{c}}$ that depends only on $x_{\mu}$ (the $\mu$-th component of the position vector). When this symmetry is not spontaneously broken, Wilson loops obey the area law and color charges transforming in the fundamental representation are confined.
One may again consider a simplified model, requiring the scalar fields to obey the fixed-length constraint ${\rm Tr}\,\Phi_{\bm{x}}^{t}\Phi_{\bm{x}}=2$. Formally, this model can be obtained by taking the limit $u,r\to\infty$ keeping the ratio $r/u=-2$ fixed. The corresponding lattice Hamiltonian reads
@@
The critical properties of the scalar fields can be monitored by using the correlation function $G_{P}$ of the gauge-invariant bilinear operator
@@ 9.4.2 Higgs phases and phase transitions
For $N_{f}\geq 2$ the lattice models show different Higgs phases associated with different symmetry-breaking patterns. The symmetries of the Higgs phases can be inferred by means of a mean-field analysis [97, 289, 99]—critical fluctuations are only relevant along the transition lines. The symmetries of the Higgs phases are indeed the symmetries of the minima of the local scalar potential $V(\Phi)={v\over 4}\,{\rm Tr}\,(\Phi^{t}\Phi)^{2}$ in the fixed-length limit ${\rm Tr}\,\Phi^{t}\Phi=2$. In the following we summarize the main properties of these phases, which depend on the number of colors $N_{c}$, of flavors $N_{f}$, and on the parameter $v$ [97, 289, 99]. Moreover, their nature also depends on the behavior of the gauge modes, which may undergo topological transitions due to the fluctuations of variables associated with the gauge-group center ${\mathbb{Z}}_{N_{c}}$. Numerical studies for $N_{c}=2$ and $N_{c}=3$ have been reported in Refs. [289, 99, 451].
@@
Two qualitatively different phase diagrams are found, depending on the number of colors $N_{c}$ and of flavors $N_{f}$. For $N_{f}\leq N_{c}^{2}-1$, independently of the specific value of $N_{c}$, there is a single Higgs phase that lies in the half-space $v<0$ for large values of $\beta$. Indeed, the mean-field analysis shows that for $\beta\to\infty$ the model orders only for $v<0$. For $v>0$, the system is disordered for $\beta\to\infty$.
For any $N_{f}$ and $N_{c}$ satisfying $N_{f}>N_{c}^{2}-1$, instead, there are two different low-temperature Higgs phases [97, 289, 99]. For $\beta\to\infty$, they correspond to the minima of the lattice potential for $v<0$ and $v>0$, respectively, and thus we again refer to these two phases as the negative-$v$ and positive-$v$ phases, respectively, although, this characterization does not hold for finite $\beta$. The two Higgs phases have different global symmetries and also different residual gauge symmetries, as defined in Sec. 3.5.3. Some details can be found in Refs. [97, 99], and in Sec. 9.4.3 for the specific case $N_{c}=2$ and $N_{f}=4$. In particular, the positive-$v$ Higgs phase has a residual ${\mathbb{Z}}_{N_{c}}$ gauge symmetry, so topological ${\mathbb{Z}}_{N_{c}}$ gauge transitions can occur when varying the inverse gauge coupling $\gamma$, as discussed in Sec. 3.5.3.
@@
The possible existence of topological transitions associated with the gauge-group center ${\mathbb{Z}}_{N_{c}}$ can also be inferred by looking at the behavior of the model in the limit $\beta\to\infty$ keeping $\kappa\equiv\beta\gamma$ fixed, see Ref. [99] for details. For $v>0$ the relevant configurations, i.e., those that minimize the scalar potential and the scalar kinetic energy $H_{K}$, are characterized by $\widetilde{U}_{{\bm{x}},\mu}=1$, so $U_{{\bm{x}},\mu}=\lambda_{{\bm{x}},\mu}\in{\mathbb{Z}}_{N_{c}}$. Therefore, in this limit the model (115) reduces to the lattice ${\mathbb{Z}}_{Q}$ gauge theory with $Q=N_{c}$ discussed in Sec. 3.5.2. In three dimensions, this lattice discrete gauge model undergoes a transition at a finite value of the gauge coupling $K$. For example,the lattice ${\mathbb{Z}}_{2}$ gauge theory [120] presents a small-$K$ confined phase and a large-$K$ deconfined phase (which may carry topological order at the quantum level [20]), separated by a critical point at $K_{c}=0.761413292(11)$ [86, 162]. If the ${\mathbb{Z}}_{N_{c}}$ gauge transition persists for finite values of $\beta$, then, when varying $\gamma$ or $\kappa$, one may have different low-temperature Higgs phases, which have the same global symmetry but differ in the large-scale behavior of the ${\mathbb{Z}}_{N_{c}}$ modes.
@@
These arguments indicate that there are two different positive-$v$ Higgs phases, which differ for the behavior of the topological modes associated with the gauge-group center ${\mathbb{Z}}_{N_{c}}$ [97, 289]. Some numerical evidence of the existence of topological ${\mathbb{Z}}_{2}$ transitions in the SU(2) gauge LNAH models with four adjoint scalar fields is presented in Ref. [99].